OpenStudy (anonymous):

Dave receives a salary of $200 a week plus a commission of 10% of his weekly sales. An equation y-mx+b represents Dave’s weekly earnings. The y-intercept is Dave’s base salary. The slope of the line is his commission. Write an equation representing Dave’s weekly earnings. If Dave sells $1500 of goods for one week, what is his salary for the week? how do i solve this?

OpenStudy (anonymous):

Dave receives a salary of $200 a week plus a commission of 10% of his weekly sales. An equation y=mx+b represents Dave's weekly earnings. The y-intercept is Dave's base salary. The slope of the line is his commission. -Write an equations representing Dave's weekly earnings. If Dave sells $1500 of goods for one week, what is his salary for the week? A. $13,00 B. $350 C.$1520 D. $130 y = .10x + 200 y = .10(1500) + 200 y = 150 + 200 y = 350 ANS : 350

OpenStudy (anonymous):

Can you fan me?

OpenStudy (anonymous):

idk how to fan you lol

OpenStudy (anonymous):

@taylorbug1225 put your mouse over my picture and there is a button that says "become a fan"